# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ 2.3 Anisotropic Global Law
We think of $\tilde{\mathcal{S}}$ as a correction term, which measures the deviation of the superoperator $\mathcal{S}$ from the “full” superoperator $M\mapsto\mathbb{E}\left[(L-\mathbb{E}L)M(L-\mathbb{E}L)\right]$, that should be vanishing in $\ell$. We also introduce
@@
The fourth and final comparison, (9d), simply states that the solution to (5) should be a good approximation for (2) for small $\tau$. For a fixed dimension $\ell\in\mathbb{N}$, it follows from the construction of $M$ using the limit point of the normal family $\{M_{1,1}^{(\tau)}\,:\,\tau>0\}$ as $\tau\to 0$ that $\|M^{(\tau)}(z)-M(z)\|\to 0$ as $\tau\to 0$. However, because we are taking $\ell\to\infty$ and $\tau\to 0$ concurently, we rely on (C2) to control this difference.
@@ 2.4 Empirical Test Error of Random Features Ridge Regression
Consider a supervised training problem with a labelled dataset $\mathcal{D}=\{(x_{j},y_{j})\}_{j=1}^{n_{\text{train}}}$ with $x_{j}\in\mathbb{R}^{n_{0}}$ and $y_{j}\in\mathbb{R}$ for every $j\in\{1,2,\ldots,n_{\text{train}}\}$. For conciseness, let $X\in\mathbb{R}^{n_{n_{\text{train}}}\times n_{0}}$ be the matrix with $j$th rows corresponding to $x_{j}^{T}$ and $y$ be the vectors of labels. We wish to learn a relation between the inputs $x_{j}$ and the outputs $y_{j}$ by fitting the parametric function $w\mapsto n^{-\frac{1}{2}}\sigma(x^{T}W)w$ for some random matrix $W\in\mathbb{R}^{n_{0}\times d}$, a $\lambda_{\sigma}$-Lipschitz activation function $\sigma:\mathbb{R}\mapsto\mathbb{R}$ applied entrywise and some weights $w\in\mathbb{R}^{d}$. This corresponds to the random features model of [55]. This model can be viewed as a two-layer neural network, where the first layer is frozen at random initialization, and only the second layer is trained. Following the setup of [46], we will assume that $W=\varphi(Z)$ for some $Z\in\mathbb{R}^{n_{0}\times d}$ with independent standard normal entries and $\varphi:\mathbb{R}\mapsto\mathbb{R}$ a $\lambda_{\varphi}$-Lipschitz function applied entrywise. The Lipschitz constants $\lambda_{\sigma}$ and $\lambda_{\varphi}$ should be independent of the dimension of the problem in the sense that, as $n\to\infty$ with $n_{\text{train}},n_{0},d\asymp n$, $\limsup_{n\to\infty}(\lambda_{\varphi}\vee\lambda_{\sigma})<\infty$.
@@
where $A=n^{-1/2}\sigma(XW)\in\mathbb{R}^{n_{\text{train}}\times d}$ denotes the random features matrix and $\delta\in\mathbb{R}_{>0}$ is the ridge parameter. In other words, we are fitting a random features model using ridge regression. The minimization problem in (10) is strongly convex and admits the closed form solution $w_{\text{ridge}}=A^{T}(AA^{T}+\delta I_{n_{\text{train}}})^{-1}y$ which is called the ridge estimator.
@@
$E_{\text{test}}:=\|\hat{y}-\hat{A}w_{\text{ridge}}\|^{2}=\|\hat{y}-\hat{A}A^{T}(AA^{T}+\delta I_{n_{\text{train}}})^{-1}y\|^{2}$ (11)
with $\hat{A}=n^{-\frac{1}{2}}\sigma(\hat{X}W)\in\mathbb{R}^{n_{\text{test}}\times d}$. This measures the performance of the model $x\mapsto n^{-\frac{1}{2}}\sigma(x^{T}W)w_{\text{ridge}}$ on $\hat{\mathcal{D}}$. If $\hat{\mathcal{D}}=\mathcal{D}$, then (11) corresponds to the training error.
@@
$\mathbb{E}[(a^{T}_{1},\hat{a}^{T}_{1})^{T}(a^{T}_{1},\hat{a}^{T}_{1})]=\begin{bmatrix}K_{AA^{T}}&K_{A\hat{A}^{T}}\\ K_{\hat{A}A^{T}}&K_{\hat{A}\hat{A}^{T}}\end{bmatrix}$
where $\{(a^{T}_{j},\hat{a}^{T}_{j})^{T}\}_{j=1}^{d}$ represent the i.i.d. columns of $A$ and $\hat{A}$ respectively. Indeed, $K_{AA^{T}}$, $K_{A\hat{A}^{T}}$, $K_{\hat{A}A^{T}}$, and $K_{\hat{A}\hat{A}^{T}}$ encode the covariance between the entries of $A$ and $\hat{A}$. Using the tools developed above, we verify [46, Conjecture 1] under an additional boundedness assumption.
@@
$\alpha=-(1+\mathrm{tr}(K_{AA^{T}}(\delta I_{n_{\text{train}}}-d\alpha K_{AA^{T}})^{-1}))^{-1}\in\mathbb{R}_{\leq 0}$
and denote $M=(\delta I_{n_{\text{train}}}-d\alpha K_{AA^{T}})^{-1}$ as well as
$\beta=\frac{\alpha^{2}\mathrm{tr}(K_{\hat{A}\hat{A}^{T}}+d\alpha K_{\hat{A}A^{T}}M(I_{n_{\text{train}}}+\delta M)K_{A\hat{A}^{T}})}{1-\|\sqrt{d}\alpha K_{AA^{T}}^{\frac{1}{2}}MK_{AA^{T}}^{\frac{1}{2}}\|_{F}^{2}}\in\mathbb{R}_{\geq 0}.$
Then, $d\beta\|K_{AA^{T}}^{\frac{1}{2}}My\|^{2}+\|d\alpha K_{\hat{A}A^{T}}My+\hat{y}\|^{2}-E_{\text{test}}\to 0$ almost surely as $n\to\infty$.
@@
To do so, we begin by expanding the squared norm in the empirical test error $E_{\text{test}}$. This expansion reveals that it can be written as a sum of bilinear forms involving $\hat{A}A^{T}(AA^{T}+\delta I_{n_{\text{train}}})^{-1}$ and its square. We recognize this term as a block of a pseudo-resolvent corresponding to the linearization in Equation 22, with the spectral parameter set to 0. This reduces the problem to an analysis of the pseudo-resolvent, along with a continuity argument.
The conditions and assumptions of Theorem 2.3 are verified as follows. First, Condition (C3) holds immediately, as the matrices $X$ and $\hat{X}$ are Lipschitz function from a Gaussian source. The Lipschitz continuity of the activation functions $\varphi$ and $\sigma$, together with the boundedness of $\|X\|$ and $\|\hat{X}\|$, allow us to see the random features matrices as compositions of Lipschitz functions applied to Gaussian sources, and consequently makes it possible to apply standard concentration results. Additionally, the assumption that the first moment of $\|A\|$ and $\|\hat{A}\|$ remains bounded as the dimension grows provides control over higher moments. This control makes it straightforward to verify that Condition (C1) holds. The Lipschitz condition in Theorem 2.3 is similarly satisfied due to the Lipschitzness of the activation functions and the boundedness of the data matrices. The stability condition, Condition (C2), is shown to hold for all $z$ in a neighborhood of the origin, based on a bound on the solution to the RDEL in operator norm, independent of the imaginary part of the spectral parameter. This analysis is case-specific, and depends on the structure of the particular linearization $L$.
@@ 2.4.1 Boundedness Assumptions
The conditions $\limsup_{n\to\infty}\mathbb{E}[\|A\|]<\infty$ and $\limsup_{n\to\infty}\mathbb{E}[\|\hat{A}\|]<\infty$ are satisfied when the data matrices exhibit approximate orthogonality, as discussed in [27, 61]. The boundedness conditions are also met with concentrated random vectors, as outlined in [45, Assumption 2]. Notably, these assumptions include the common case of i.i.d. standard normal entries in independent data matrices, a widely studied scenario in the literature.
It is worth noting that under the assumption that the norms of the training and test random features matrices are bounded in expectation, [46] conjectured Theorem 2.4 and provided a sketch of the proof. Moreover, [45] proved a similar result in the context of random Fourier features. While the conclusion of Theorem 2.4 may seem unsurprising, to the best of our knowledge, this is the first time that the theorem has been rigorously established at this level of generality. Beyond being a significant theoretical result, the derivation of Theorem 2.4 also demonstrates the broader utility of our framework based on the Dyson equation for correlated linearizations to solve problems at the interface of random matrix theory and machine learning.
@@ 2.4.2 Bounded Denominator
While not obvious at first, we show in Lemma 6.7 that $1-\|\sqrt{d}\alpha K_{AA^{T}}^{1/2}MK_{AA^{T}}^{1/2}\|_{F}^{2}$ is positive and bounded away from 00 as $n\to\infty$ in the setting of Theorem 2.4. This implies that $\beta$, and therefore $E_{\text{test}}$, is well-behaved in the proportional limit.
@@ 2.4.3 Data Assumptions and Real-World Relevance
Our assumptions concerning the norms of $A$ and $\hat{A}$ implicitly impose constraints on the data matrices $X$ and $\hat{X}$. However, conditioning on these matrices allows us to establish asymptotic equivalence without requiring restrictive distributional assumptions. Consequently, our results is applicable to a broad spectrum of data matrices, offering a more accurate model for real-world datasets. For instance, as illustrated in Figure 2, we observe a striking alignment between the empirical test error $E_{\text{test}}$ and the deterministic approximation provided by Theorem 2.4 across various dimensions and ridge parameters when the data is sourced from real-world flattened image classification datasets such as MNIST [22], Fashion-MNIST [63], CIFAR-10 [43], and CIFAR-100 [43]. Notably, the agreement between empirical simulations and the theoretical prediction of Theorem 2.4 holds even for datasets with anisotropic features.
@@ 2.4.4 Extension to Deep Random Features
$x\in\mathbb{R}^{n_{0}}\mapsto n_{k}^{-\frac{1}{2}}\sigma\left(n^{-\frac{1}{2}}_{k-1}\sigma\left(\cdots n^{-\frac{1}{2}}\sigma\left(n^{-\frac{1}{2}}_{1}\sigma(x^{T}W_{1})W_{2}\right)\right)W_{k}\right)w\in\mathbb{R}^{n_{k}}.$
Here, $k\in\mathbb{N}$ corresponds to the number of layers and $\{W_{j}\in\mathbb{R}^{n_{j-1}\times n_{k}}\}_{j=1}^{k}$ is a collection of random matrices with independent standard normal entries. Under the assumptions of [27], it follows from [27, Lemma D.4] and the equivalent characterizations of sub-exponential random variables [60, Proposition 2.7.1] that the norm of the conjugate kernel matrices $AA^{T}$ and $\hat{A}\hat{A}^{T}$ have bounded moments. This allows for a direct extension of Theorem 2.4 to deep random features.
Figure 2: $E_{\text{test}}$ vs the deterministic approximation given in Theorem 2.4 for various flattened image classification datasets with different sizes of hidden layers $d$ and ridge parameter $\delta$. Sine activation ($\sigma=\sin$), $n_{\text{train}}=1500$, $n_{\text{test}}=1000$. Upper left: MNIST [22]; Upper right: Fashion-MNIST [63]; Lower left: CIFAR-10 [43]; Lower right: CIFAR-100 [43].

### v4

@@ 2.3 Anisotropic Global Law
We think of $\tilde{\mathcal{S}}$ as a correction term, which measures the deviation of the superoperator $\mathcal{S}$ from the “full” superoperator $M\mapsto\mathbb{E}\left[(L-\mathbb{E}L)M(L-\mathbb{E}L)\right]$, that should be vanishing for $\ell\to\infty$. We also introduce
@@
The fourth and final comparison, (9d), simply states that the solution to (5) should be a good approximation for (2) for small $\tau$. For a fixed dimension $\ell\in\mathbb{N}$, it follows from the construction of $M$ using the limit point of the normal family $\{M_{1,1}^{(\tau)}\,:\,\tau>0\}$ as $\tau\to 0$ that $\|M^{(\tau)}(z)-M(z)\|\to 0$ as $\tau\to 0$. However, because we are taking $\ell\to\infty$ and $\tau\to 0$ concurrently, we rely on (C2) to control this difference.
@@ 2.4 Empirical Test Error of Random Features Ridge Regression
Consider a supervised training problem with a labelled dataset $\mathcal{D}=\{(x_{j},y_{j})\}_{j=1}^{n_{\text{train}}}$ with $x_{j}\in\mathbb{R}^{n_{0}}$ and $y_{j}\in\mathbb{R}$ for every $j\in\{1,2,\ldots,n_{\text{train}}\}$. For conciseness, let $X\in\mathbb{R}^{n_{n_{\text{train}}}\times n_{0}}$ be the matrix with $j$th rows corresponding to $x_{j}^{\top}$ and $y$ be the vectors of labels. We wish to learn a relation between the inputs $x_{j}$ and the outputs $y_{j}$ by fitting the parametric function $w\mapsto n^{-\frac{1}{2}}\sigma(x^{\top}W)w$ for some random matrix $W\in\mathbb{R}^{n_{0}\times d}$, a $\lambda_{\sigma}$-Lipschitz activation function $\sigma:\mathbb{R}\mapsto\mathbb{R}$ applied entrywise and some weights $w\in\mathbb{R}^{d}$. This corresponds to the random features model of [Rahimi_RF]. This model can be viewed as a two-layer neural network, where the first layer is frozen at random initialization, and only the second layer is trained. Following the setup of [louart_random-matrix-approach], we will assume that $W=\varphi(Z)$ for some $Z\in\mathbb{R}^{n_{0}\times d}$ with independent standard normal entries and $\varphi:\mathbb{R}\mapsto\mathbb{R}$ a $\lambda_{\varphi}$-Lipschitz function applied entrywise. The Lipschitz constants $\lambda_{\sigma}$ and $\lambda_{\varphi}$ should be independent of the dimension of the problem in the sense that, as $n\to\infty$ with $n_{\text{train}},n_{0},d\asymp n$, $\limsup_{n\to\infty}(\lambda_{\varphi}\vee\lambda_{\sigma})<\infty$.
@@
where $A=n^{-1/2}\sigma(XW)\in\mathbb{R}^{n_{\text{train}}\times d}$ denotes the random features matrix and $\delta\in\mathbb{R}_{>0}$ is the ridge parameter. In other words, we are fitting a random features model using ridge regression. The minimization problem in (10) is strongly convex and admits the closed form solution $w_{\text{ridge}}=A^{\top}(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}y$ which is called the ridge estimator.
@@
$E_{\text{test}}:=\|\hat{y}-\hat{A}w_{\text{ridge}}\|^{2}=\|\hat{y}-\hat{A}A^{\top}(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}y\|^{2}$ (11)
with $\hat{A}=n^{-\frac{1}{2}}\sigma(\hat{X}W)\in\mathbb{R}^{n_{\text{test}}\times d}$. This measures the performance of the model $x\mapsto n^{-\frac{1}{2}}\sigma(x^{\top}W)w_{\text{ridge}}$ on $\hat{\mathcal{D}}$. If $\hat{\mathcal{D}}=\mathcal{D}$, then (11) corresponds to the training error.
@@
$\mathbb{E}[(a^{\top}_{1},\hat{a}^{\top}_{1})^{\top}(a^{\top}_{1},\hat{a}^{\top}_{1})]=\begin{bmatrix}K_{AA^{\top}}&K_{A\hat{A}^{\top}}\\ K_{\hat{A}A^{\top}}&K_{\hat{A}\hat{A}^{\top}}\end{bmatrix}$
where $\{(a^{\top}_{j},\hat{a}^{\top}_{j})^{\top}\}_{j=1}^{d}$ represent the i.i.d. columns of $A$ and $\hat{A}$ respectively. Indeed, $K_{AA^{\top}}$, $K_{A\hat{A}^{\top}}$, $K_{\hat{A}A^{\top}}$, and $K_{\hat{A}\hat{A}^{\top}}$ encode the covariance between the entries of $A$ and $\hat{A}$. Using the tools developed above, we verify [louart_random-matrix-approach, Conjecture 1] under an additional boundedness assumption.
@@
$\alpha=-(1+\mathrm{tr}(K_{AA^{\top}}(\delta I_{n_{\text{train}}}-d\alpha K_{AA^{\top}})^{-1}))^{-1}\in\mathbb{R}_{\leq 0}$
and denote $M=(\delta I_{n_{\text{train}}}-d\alpha K_{AA^{\top}})^{-1}$ as well as
$\beta=\frac{\alpha^{2}\mathrm{tr}(K_{\hat{A}\hat{A}^{\top}}+d\alpha K_{\hat{A}A^{\top}}M(I_{n_{\text{train}}}+\delta M)K_{A\hat{A}^{\top}})}{1-\|\sqrt{d}\alpha K_{AA^{\top}}^{\frac{1}{2}}MK_{AA^{\top}}^{\frac{1}{2}}\|_{F}^{2}}\in\mathbb{R}_{\geq 0}.$
Then, $d\beta\|K_{AA^{\top}}^{\frac{1}{2}}My\|^{2}+\|d\alpha K_{\hat{A}A^{\top}}My+\hat{y}\|^{2}-E_{\text{test}}\to 0$ almost surely as $n\to\infty$.
@@
To do so, we begin by expanding the squared norm in the empirical test error $E_{\text{test}}$. This expansion reveals that it can be written as a sum of bilinear forms involving $\hat{A}A^{\top}(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}$ and its square. We recognize this term as a block of a pseudo-resolvent corresponding to the linearization in Equation 22, with the spectral parameter set to 0. This reduces the problem to an analysis of the pseudo-resolvent, along with a continuity argument.
The conditions and assumptions of Theorem 2.2 are verified as follows. First, Condition (C3) holds immediately, as the matrices $X$ and $\hat{X}$ are Lipschitz functions from a Gaussian source. The Lipschitz continuity of the activation functions $\varphi$ and $\sigma$, together with the boundedness of $\|X\|$ and $\|\hat{X}\|$, allow us to see the random features matrices as compositions of Lipschitz functions applied to Gaussian sources, and consequently makes it possible to apply standard concentration results. Additionally, the assumption that the first moment of $\|A\|$ and $\|\hat{A}\|$ remains bounded as the dimension grows provides control over higher moments. This control makes it straightforward to verify that Condition (C1) holds. The Lipschitz condition in Theorem 2.2 is similarly satisfied due to the Lipschitzness of the activation functions and the boundedness of the data matrices. The stability condition, Condition (C2), is shown to hold for all $z$ in a neighborhood of the origin, based on a bound on the solution to the RDEL in operator norm, independent of the imaginary part of the spectral parameter. This analysis is case-specific, and depends on the structure of the particular linearization $L$.
@@
Let us briefly comment on the interpretation of Theorem 2.3. On the one hand, it shows that the empirical test error $E_{\text{test}}$ concentrates around a deterministic quantity that depends only on the first-order and second-order statistics of the random features matrices $A$ and $\hat{A}$. This establishes a Gaussian equivalence principle for the empirical test error of random features ridge regression, as discussed in Section 2.4.5. The Gaussian equivalence principle reduces the analysis of random features ridge regression to that of a surrogate Gaussian model, thereby simplifying subsequent analyses of this model. Furthermore, the Gaussian universality of the test error implies that, once the relevant covariance operators are fixed, the distributional details of the random features beyond their second moments become immaterial. This can be viewed as a fundamental limit on statistical power: only the Gaussian approximation of the post-activation features matters for the quality of the fit.
On the other hand, because the deterministic limit is available in closed form, it enables a systematic study of how model parameters such as the activation function $\sigma$, the ridge parameter $\delta$, or the mismatch between training and test distributions (encoded in the covariance matrices) affect the asymptotic performance. For instance, Section 2.4.7 leverages the explicit formula in Theorem 2.3 to link the asymptotic performance of random features ridge regression to that of kernel ridge regression, where the effective regularization is implicitly increased by the random features model.
@@ 2.4.1 Boundedness Assumptions
The conditions
$\limsup_{n\to\infty}\mathbb{E}[\|A\|]<\infty\quad\text{and}\quad\limsup_{n\to\infty}\mathbb{E}[\|\hat{A}\|$$\infty$
are satisfied when the data matrices exhibit approximate orthogonality, as discussed in [fan_ck, wang2024nonlinear]. The boundedness conditions are also met with concentrated random vectors, as outlined in [Liao_2021, Assumption 2]. Notably, these assumptions include the common case of i.i.d. standard normal entries in independent data matrices, a widely studied scenario in the literature.
It is worth noting that under the assumption that the norms of the training and test random features matrices are bounded in expectation, [louart_random-matrix-approach] conjectured Theorem 2.3 and provided a sketch of the proof. Moreover, [Liao_2021] proved a similar result in the context of random Fourier features. While the conclusion of Theorem 2.3 may seem unsurprising, to the best of our knowledge, this is the first time that the theorem has been rigorously established at this level of generality. Beyond being a significant theoretical result, the derivation of Theorem 2.3 also demonstrates the broader utility of our framework based on the Dyson equation for correlated linearizations to solve problems at the interface of random matrix theory and machine learning.
@@ 2.4.2 Bounded Denominator
While not obvious at first, we show in Lemma 6.7 that $1-\|\sqrt{d}\alpha K_{AA^{\top}}^{1/2}MK_{AA^{\top}}^{1/2}\|_{F}^{2}$ is positive and bounded away from 0 as $n\to\infty$ in the setting of Theorem 2.3. This implies that $\beta$, and therefore $E_{\text{test}}$, is well-behaved in the proportional limit.
@@ 2.4.3 Data Assumptions and Real-World Relevance
Our assumptions concerning the norms of $A$ and $\hat{A}$ implicitly impose constraints on the data matrices $X$ and $\hat{X}$. However, conditioning on these matrices allows us to establish asymptotic equivalence without requiring restrictive distributional assumptions. Consequently, our results are applicable to a broad spectrum of data matrices, offering a more accurate model for real-world datasets. For instance, as illustrated in Figure 2, we observe a striking alignment between the empirical test error $E_{\text{test}}$ and the deterministic approximation provided by Theorem 2.3 across various dimensions and ridge parameters when the data is sourced from real-world flattened image classification datasets such as MNIST [deng2012mnist], Fashion-MNIST [xiao2017fashionmnist], CIFAR-10 [Krizhevsky09learningmultiple], and CIFAR-100 [Krizhevsky09learningmultiple]. Notably, the agreement between empirical simulations and the theoretical prediction of Theorem 2.3 holds even for datasets with anisotropic features.
Figure 2: $E_{\text{test}}$ vs the deterministic approximation given in Theorem 2.3 for various flattened image classification datasets with different sizes of hidden layers $d$ and ridge parameter $\delta$. Sine activation ($\sigma=\sin$), $n_{\text{train}}=1500$, $n_{\text{test}}=1000$. Upper left: MNIST [deng2012mnist]; Upper right: Fashion-MNIST [xiao2017fashionmnist]; Lower left: CIFAR-10 [Krizhevsky09learningmultiple]; Lower right: CIFAR-100 [Krizhevsky09learningmultiple].
@@ 2.4.4 Extension to Deep Random Features
$x\in\mathbb{R}^{n_{0}}\mapsto n_{k}^{-\frac{1}{2}}\sigma\left(n^{-\frac{1}{2}}_{k-1}\sigma\left(\cdots n^{-\frac{1}{2}}_{2}\sigma\left(n^{-\frac{1}{2}}_{1}\sigma(x^{\top}W_{1})W_{2}\right)\right)W_{k}\right)w\in\mathbb{R}^{n_{k}}.$
Here, $k\in\mathbb{N}$ corresponds to the number of layers and $\{W_{j}\in\mathbb{R}^{n_{j-1}\times n_{k}}\}_{j=1}^{k}$ is a collection of random matrices with independent standard normal entries.
More precisely, suppose we adopt the setting of [fan_ck], which considers deep random features models in the proportional regime where $n_{1}\asymp n_{2}\asymp\cdots\asymp n_{k}\asymp n$ with standard normal weights, an activation function $\sigma$ that is Lipschitz with some additional regularity assumptions, and input data matrices $X$ satisfying the notion of approximate pairwise orthogonality from [fan_ck, Definition 3.1]. This notion notably ensures that $\|X\|$ and $\mathbb{E}[\|n^{-1/2}\sigma(XW)\|]$ remain bounded as $n,d\to\infty$ for $W$ with i.i.d. $\mathcal{N}(0,1)$ entries [fan_ck, Definition 3.1 and Lemma D.4]. Moreover, [fan_ck, Lemma D.1] shows that approximate orthogonality propagates through the layers with overwhelming probability as $n,d\to\infty$.
Hence, let both $X$ and $\hat{X}$ satisfy these conditions. Define
$\tilde{X}=n_{k-1}^{-\frac{1}{2}}\sigma(\cdots n_{1}^{-\frac{1}{2}}\sigma(n_{0}^{-\frac{1}{2}}XW_{1})\cdots W_{k-1})\in\mathbb{R}^{n_{\text{train}}\times n_{k-1}}$
and similarly for $\hat{\tilde{X}}$ with $\hat{X}$ in place of $X$. Let $W:=W_{k}$ and define the random features matrices at the last layer
$A=n^{-1/2}\sigma(\tilde{X}W)\in\mathbb{R}^{n_{\text{train}}\times n_{k}},\qquad\hat{A}=n^{-1/2}\sigma(\hat{\tilde{X}}W)\in\mathbb{R}^{n_{\text{test}}\times n_{k}}.$
Conditioning on an event of overwhelming probability involving the weights $\{W_{j}\}_{j=1}^{k-1}$ and using the approximate orthogonality of $X$ and $\hat{X}$, it follows from [fan_ck, Lemma D.1] that the conditions of Theorem 2.3 are satisfied. Applying Theorem 2.3 then yields an object that does not depend on the last-layer weights $W$ and is asymptotically equivalent to the empirical test error of the deep random features model with overwhelming probability as $n,d\to\infty$. Admittedly, this object still depends on the earlier weights $\{W_{j}\}_{j=1}^{k-1}$ through $\tilde{X}$ and $\hat{\tilde{X}}$, and further analysis is required to obtain a fully explicit deterministic equivalent, but this is beyond the scope of this paper.